OpenStudy (fernandez555):

Which of the following represents the area of a rectangle whose length is 3x + 5 and whose width is x - 2? A. 3x2 - x - 10 B. 3x2 - 10 C. 3x2 + x - 10 D. 3x2 - 11x - 10

jimthompson5910 (jim_thompson5910):

A = L*W A = (3x+5)*(x-2) A = 3x(x-2) + 5(x-2) A = ???

OpenStudy (fernandez555):

@jim_thompson5910 3x2 + x - 10 ???

jimthompson5910 (jim_thompson5910):

What do you get when you distribute out 3x(x-2) + 5(x-2)

OpenStudy (fernandez555):

3x2 and 10 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

3x(x-2) + 5(x-2) 3x*x+3x*(-2) + 5(x) + 5(-2) 3x^2 - 6x + 5x - 10 last step is???

OpenStudy (fernandez555):

-x + 3x^2 -1??? @jim_thompson5910

OpenStudy (fernandez555):

@jim_thompson5910 wait i simplified. ahhhhh idk, im lost i havent done this in forever

jimthompson5910 (jim_thompson5910):

3x(x-2) + 5(x-2) 3x*x+3x*(-2) + 5(x) + 5(-2) 3x^2 - 6x + 5x - 10 3x^2 - x - 10

jimthompson5910 (jim_thompson5910):

Not sure where the -1 came from